OpenStudy (anonymous):

integral from 0 to 1 of cos(arcsinx)

OpenStudy (anonymous):

Let \[\theta =\sin^{-1} x\]

OpenStudy (anonymous):

Therefore \[x/1=\sin \theta =opp/hyp\]Draw my triangle x is opp, 1 hyp, sq rt (1-x^2) adj

OpenStudy (anonymous):

From that triangle you have info to find cos theta which you would integrate

OpenStudy (anonymous):

do you integrate cos theta or cos (sqrt(1-x^2)) ?

OpenStudy (anonymous):

cos theta in and of itself does not contain the info in your problem. However cos theta (which essentially is your problem) is equal to sq root (1-x^2) which contains all relevant info. So integrate sq rt (1-x^2).

OpenStudy (anonymous):

how do you integrate sq rt(1-x^2)?

OpenStudy (anonymous):

You have to find relevant trig sub.

OpenStudy (anonymous):

Your instructor gives very interesting problems. What school? what course?

OpenStudy (anonymous):

University of Miami: Calculus II

OpenStudy (anonymous):

Go 'Canes

OpenStudy (anonymous):

Lol! Thanks